8 October 2013: Tuesday

LDAP:

  • Update:
    • One issue from the update.  A new employee that did not find a match.  A foreign student who does not have an SSN.  Matched and merged.

NetWorker:

  • New Hardware:
    • The staging of all valid data from UVM.SATABeast.001 completed at 9:16 PM last night.  I have unmounted the volume, and removed both the device and the empty volume from the media database.
    • Starting the process of emptying UVM.SATABeast.002.  That one has 6,777 save sets on it.
  • COMIS groups:
    • Need to examine these groups and rebalance them for their weekends.  Have finally figured out why the MED23B group runs over so much.  It is dealing with over 16 TB of data and trying to push that through a 1Gb/s NIC and I’m amazed that it is capable of doing so in 64hours.  However, the goal is 48 hours, so that has to be fixed.
    • Analysis shows:
      • First Friday: 105 Save sets, 2TB, 3.5M files, 7.9H (8/2 11PM – 8/3 6:58AM)
      • Second Friday: 5 Save sets, 8TB, 6.6M files, 40.75H (8/9 11PM – 8/11 3:44PM)
      • Third Friday: 10 Save sets, 16.5TB, 3.8M files, 63.5H (8/16 11PM – 8/19 3:27PM)
      • Fourth Friday: 12 Save Sets, 13.6TB, 13.3M files, 63.75H (8/30 11PM – 9/2 3:45PM)
  • Old Hardware:
    • The replacement disk drive arrived and stornode1 is back to fully protected state.
  • Tape Dancing:
    • 97 tapes moved over to the vaults.
    • 64 tapes brought back.
    • Started relabeling the 64.

Tags:

Comments are closed.